The total value of two bank accounts is $1100. The annual interest earned on the checking account is 5%, while the CD pays 6%. The total interest earned is $61. How much was deposited into each account?
Answer by stanbon(75887)

The total value of two bank accounts is $1100. The annual interest earned on the checking account is 5%, while the CD pays 6%. The total interest earned is $61. How much was deposited into each account?
---------------------
Quantity Equation:: x + y = 1100
Interest Equation:: 5x + 6y = 6100
-----------------------------------
Multiply thru 1st equation by 5:
5x + 5y = 5500
5x + 6y = 6100
----
Solve for "y":
y = $600 (amt. invested at 6%)
-------
Solve for "x":
x + 600 = 1100
x = $500 (amt. invested at 5%)
====================
Cheers,
Stan H.
===========